# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ Appendix B Pre-training
$\mathcal{H}(\hat{\boldsymbol{E}},\hat{\boldsymbol{D}})=\widetilde{\mathcal{O}}\left(\mathcal{M}^{-\frac{1}{d+2}}+\delta_{0}\right).$
The proof is complete.
Appendix C Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$
In subsection C.1, we prove Lemma 4.4, which gives an upper bound for the approximation error. This proof is structured into two parts. In subsection C.2, we prove Lemma 4.5, which provides an upper bound for the statistical error. In subsection C.3, we prove Theorem 4.6. In the last subsection C.4, we prove Theorem 4.7.
C.1 Approximation Error
Approximation on $[0,T]\times\mathcal{K}$.
We choose $\mathcal{K}=\{\mathbf{x}|\|\mathbf{x}\|_{\infty}\leq r\}$, where $r>0$ will be determined later. On $[0,T]\times\mathcal{K}$, we approximate $i-$coordinate maps of $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$, and then obtain an approximation of $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ by concatenation.
We first rescale the input by $\mathbf{x}^{\prime}=\frac{1}{2r}(\mathbf{x}+r\mathbf{1})$ and $t^{\prime}=\frac{t}{T}$, so that the transformed space is $[0,1]\times[0,1]^{d^{*}}$. We defined the rescaled function as $\mathbf{Q}(t^{\prime},\mathbf{x}^{\prime})=\nabla\log{q}(\sqrt{1-Tt^{\prime}}\sigma,2r\mathbf{x}^{\prime}-r\mathbf{1})$, then $\mathbf{Q}(t^{\prime},\mathbf{x}^{\prime})$ is $2\xi r$-Lipschitz in $\mathbf{x}^{\prime}$. We denote
$\tau(r):=\mathop{\rm{sup}}_{t\in[0,T]}\mathop{\rm{sup}}_{\mathbf{x}\in[-r,r]^{d^{*}}}\|\partial_{t}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})\|.$
By Lemma A.4, we have $\tau(r)=\mathcal{O}\left(\frac{{d^{*}}^{3/2}(r+1)}{(1-T)^{3}}\right)$, then $\mathbf{Q}(t^{\prime},\mathbf{x}^{\prime})$ is $T\tau(r)-$Lipschitz over $t^{\prime}$. Now our goal is to approximate $\mathbf{Q}(t^{\prime},\mathbf{x}^{\prime})$ on $[0,1]\times[0,1]^{d^{*}}$.
Second, we partition the time interval $[0,1]$ into non-overlapping sub-intervals of length $e_{1}$. We also partition $[0,1]^{d^{*}}$ into non-overlapping hypercubes with equal edge length $e_{2}$. $e_{1}$ and $e_{2}$ will be chosen depending on the desired approximation error. We denote $N_{1}=\lceil\frac{1}{e_{1}}\rceil$ and $N_{2}=\lceil\frac{1}{e_{2}}\rceil$.
Let $\mathbf{m}=[m_{1},m_{2},\cdots,m_{d^{*}}]^{\top}\in[N_{2}]^{d^{*}}$ be a multi-index. We define $\overline{\mathbf{Q}}(t^{\prime},\mathbf{x}^{\prime})$ as
$\overline{Q}_{i}(t^{\prime},\mathbf{x}^{\prime}):=\sum_{j\in[N_{1}],\mathbf{m}\in[N_{2}]^{d^{*}}}Q_{i}\left(\frac{j}{N_{1}},\frac{\mathbf{m}}{N_{2}}\right)\Psi_{j,\mathbf{m}}(t^{\prime},\mathbf{x}^{\prime}),i=1,2,\cdots,d^{*},$
where $\Psi_{j,\mathbf{m}}(t^{\prime},\mathbf{x}^{\prime})$ is a partition of unity function which satisfies that
$\sum_{j\in[N_{1}],\mathbf{m}\in[N_{2}]^{d^{*}}}\Psi_{j,\mathbf{m}}(t^{\prime},\mathbf{x}^{\prime})\equiv 1$
on $[0,1]\times[0,1]^{d^{*}}$. We choose $\Psi_{j,\mathbf{m}}$ as a product of coordinate-wise trapezoid functions:
$\Psi_{j,\mathbf{m}}(t^{\prime},\mathbf{x}^{\prime}):=\psi\left(3N_{1}\left(t^{\prime}-\frac{j}{N_{1}}\right)\right)\prod_{i=1}^{d^{*}}\psi\left(3N_{2}\left(x_{i}^{\prime}-\frac{m_{i}}{N_{2}}\right)\right),$
where $\psi$ is a trapezoid function,
$\psi(a):=\begin{cases}1,&|a|<1,\\ 2-|a|,&|a|\in[1,2],\\ 0,&|a|>2.\end{cases}$
We claim that
  • $\overline{Q}_{i}(t^{\prime},\mathbf{x}^{\prime})$ is an approximation of $Q_{i}(t^{\prime},\mathbf{x}^{\prime})$;
  • $\overline{Q}_{i}(t^{\prime},\mathbf{x}^{\prime})$ can be implemented by a ReLU neural network $S_{i}(t^{\prime},\mathbf{x}^{\prime})$ with an approximating error small enough.
Both claims are verified in [CLZZ20, Lemma 10], where we only need to substitute the Lipschitz constant $2\xi r$ and $T\tau(r)$ into the error analysis. By concatenating $S_{i}(t^{\prime},\mathbf{x}^{\prime}),i=1,2,\cdots,d^{*}$ together, we construct $\mathbf{S}=[S_{1},S_{2},\cdots,S_{d^{*}}]^{\top}$. Given $\epsilon>0$, we have
$\mathop{\rm{sup}}_{t^{\prime},\mathbf{x}^{\prime}\in[0,1]\times[0,1]^{d^{*}}}\|\mathbf{S}(t^{\prime},\mathbf{x}^{\prime})-\mathbf{Q}(t^{\prime},\mathbf{x}^{\prime})\|_{\infty}\leq\epsilon,$
where the neural network configuration is
$L=\mathcal{O}\left(\log{\frac{1}{\epsilon}}+d^{*}\right),M=\mathcal{O}\left(d^{*}\tau(r)(\xi r)^{d^{*}}\epsilon^{-(d^{*}+1)}\right),$
$J=\mathcal{O}\left(d^{*}\tau(r)(\xi r)^{d^{*}}\epsilon^{-(d^{*}+1)}\left(\log{\frac{1}{\epsilon}+d^{*}}\right)\right),$
$K=\mathcal{O}\left(\frac{\sqrt{d^{*}}(r+1)}{1-T}\right),\kappa=\mathcal{O}\left({\rm{max}}\{1,\xi r,T\tau(r)\}\right).$
Here, we set $e_{1}=\mathcal{O}\left(\frac{\epsilon}{T\tau(r)}\right)$ and $e_{2}=\mathcal{O}\left(\frac{\epsilon}{\xi r}\right)$. The output range $K$ is computed by
$\mathop{\text{sup}}_{t\in[0,T],\mathbf{x}\in[-r,r]^{d^{*}}}\|\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})\|\leq\frac{\|\mathbf{x}\|+\|\mathbf{y}\|}{\sigma^{2}(1-T)}\leq\frac{\sqrt{d^{*}}(r+1)}{\sigma^{2}(1-T)}=:K.$
Furthermore, $\mathbf{S}(t^{\prime},\mathbf{x}^{\prime})$ is Lipschitz continuous in $\mathbf{x}^{\prime}$, i.e., for any $\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime}\in[0,1]^{d^{*}}$ and $t^{\prime}\in[0,1]$, it holds
$\|\mathbf{S}(t^{\prime},\mathbf{x}_{1}^{\prime})-\mathbf{S}(t^{\prime},\mathbf{x}_{2}^{\prime})\|_{\infty}\leq 20d^{*}\xi{r}\|{\mathbf{x}_{1}^{\prime}-\mathbf{x}_{2}^{\prime}}\|.$
$\mathbf{S}(t^{\prime},\mathbf{x}^{\prime})$ is also Lipschitz continuous in $t$, i.e., for any $t_{1},t_{2}\in[0,1]$ and $\mathbf{x}^{\prime}\in[0,1]^{d^{*}}$, it holds
$\|\mathbf{S}(t_{1}^{\prime},\mathbf{x}^{\prime})-\mathbf{S}(t_{2}^{\prime},\mathbf{x}^{\prime})\|_{\infty}\leq 10T\tau(r)|t_{1}^{\prime}-t_{2}^{\prime}|.$
We denote $\mathbf{s}(t,\mathbf{x})=\mathbf{S}\left(\frac{t}{T},\frac{1}{2r}(\mathbf{x}+r\mathbf{1})\right)$, then for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{K}$ and $t\in[0,T]$, it holds
$\|\mathbf{s}(t,\mathbf{x}_{1})-\mathbf{s}(t,\mathbf{x}_{2})\|_{\infty}\leq 10d^{*}\xi\|{\mathbf{x}_{1}-\mathbf{x}_{2}}\|.$ (11)
And for any $t_{1},t_{2}\in[0,T]$ and $\mathbf{x}\in\mathcal{K}$, it holds
$\|\mathbf{s}(t_{1},\mathbf{x})-\mathbf{s}(t_{2},\mathbf{x})\|_{\infty}\leq 10\tau(r)|t_{1}-t_{2}|.$
Global Lipschitz. In order to ensure global Lipschitz continuity of $\mathbf{s}(t,\mathbf{x})$ with respect to $\mathbf{x}$, we introduce the following lemma.
Lemma C.1.
For $\mathcal{K}=\{\mathbf{x}\in\mathbb{R}^{d^{*}}|\|{\mathbf{x}}\|_{\infty}\leq r\}$, there exists an $1$-Lipshitz map $\mathcal{T}_{\mathcal{K}}:\mathbb{R}^{d^{*}}\rightarrow\mathcal{K}$, which satisfies $\mathcal{T}_{\mathcal{K}}(\mathbf{x})=\mathbf{x}$, for any $\mathbf{x}\in\mathcal{K}$. And $\mathcal{T}_{\mathcal{K}}$ can be expressed as an 2-layer ReLU network with width order $\mathcal{O}(d^{*})$.
@@ C.1 Approximation Error
We consider the following univariate real-valued function
$f_{r}(x):=\begin{cases}r,&x>r\\ x,&x\in[-r,r]\\ -r,&x<-r\end{cases}$
and define the map $\mathcal{T}_{\mathcal{K}}(\mathbf{x})=(f_{r}(x_{1}),f_{r}(x_{2}),\cdots,f_{r}(x_{d^{*}}))^{\top}$. Then, for any $\mathbf{x}\in\mathcal{K}$, $\mathcal{T}_{\mathcal{K}}(\mathbf{x})=\mathbf{x}$, and for any $\mathbf{x}\in\mathcal{K}^{c}$, $\mathcal{T}_{\mathcal{K}}(\mathbf{x})\in\partial\mathcal{K}$. By a simple calculation, we have
$\|\mathcal{T}_{\mathcal{K}}(\mathbf{x})-\mathcal{T}_{\mathcal{K}}(\mathbf{y})\|_{\infty}$ $=\mathop{\max}_{i=1,2,\cdots,d^{*}}|f_{r}(x_{i})-f_{r}(y_{i})|$
$\leq\mathop{\max}_{i=1,2,\cdots,d^{*}}|x_{i}-y_{i}|=\|\mathbf{x}-\mathbf{y}\|_{\infty}\leq\|\mathbf{x}-\mathbf{y}\|.$
And it is easy to check
$f_{r}(x)=\mathrm{ReLU}(x)-\mathrm{ReLU}(-x)+\mathrm{ReLU}(-x-r)-\mathrm{ReLU}(x-r).$
@@
We now consider the neural network $\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}))$. It preserves the approximation capability of $\mathbf{s}(t,\mathbf{x})$, i.e.,
$\sup_{t,\mathbf{x}\in[0,T]\times\mathcal{K}}\|\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}))-\nabla\log q(\sqrt{1-t}\sigma,\mathbf{x})\|_{\infty}$
$=$ $\sup_{t,\mathbf{x}\in[0,T]\times\mathcal{K}}\|\mathbf{s}(t,\mathbf{x})-\nabla\log q(\sqrt{1-t}\sigma,\mathbf{x})\|_{\infty}\leq\epsilon.$
And it holds the upper bound, i.e.,
$\sup_{t,\mathbf{x}\in[0,T]\times\mathbb{R}^{d^{*}}}\|\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}))\|=\sup_{t,\mathbf{x}\in[0,T]\times\mathcal{K}}\|\mathbf{s}(t,\mathbf{x})\|=K.$
For any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{R}^{d^{*}}$ and $t\in[0,T]$, it holds
$\|\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}_{1}))-\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}_{2}))\|_{\infty}$ $\leq 10d^{*}\xi\|\mathcal{T}_{\mathcal{K}}(\mathbf{x}_{1})-\mathcal{T}_{\mathcal{K}}(\mathbf{x}_{2})\|$
$\leq 10d^{*}\xi\|\mathbf{x}_{1}-\mathbf{x}_{2}\|.$
Therefore, $\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}))$ is global Lipschitz continuous with respect to $\mathbf{x}$. For notational convenience, we still use $\mathbf{s}(t,\mathbf{x})$ to denote $\mathbf{s}(t,\mathcal{T}_{\mathcal{K}}(\mathbf{x}))$.
$L^{2}$ approximation error.
The $L^{2}$ approximation error can be decomposed into two terms,
$\mathbb{E}\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}$ (12)
$=\mathbb{E}\left[\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathbb{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq{r}\}}\right]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\mathbb{E}\left[\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathbb{I}_{\{\|\mathbf{x}_{t}\|_{\infty}>{r}\}}\right].$
Then, it suffices to bound the two terms on the right hand of (12). The first term satisfies
$\mathbb{E}\left[\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathbb{I}_{\{\|\mathbf{x}_{t}\|_{\infty}\leq{r}\}}\right]$
$\leq d^{*}\sup_{t\in[0,T],\mathbf{x}_{t}\in[-r,r]^{d^{*}}}\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|_{\infty}^{2}$
$\leq d^{*}\epsilon^{2}.$
The second term satisfies
$\leavevmode\nobreak\ \mathbb{E}\left[\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathbb{I}_{\{\|\mathbf{x}_{t}\|_{\infty}>{r}\}}\right]$

### v2

@@ Appendix A Pre-training
$\mathcal{H}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})=\widetilde{\mathcal{O}}\left(\mathcal{M}^{-\frac{1}{d+2}}+\delta_{0}\right).$
The proof is complete. ∎
Appendix B Approximation Error
Building on the preliminary lemmas in Sections F.1-F.2, we now proceed to the proof of Lemma 4.3 by following [OAS23, FYWC24]. Recall that the score function $\nabla\log q_{t}(\mathbf{x})$ can be rewritten as
$\nabla\log q_{t}(\mathbf{x})=\frac{\nabla q_{t}(\mathbf{x})}{q_{t}(\mathbf{x})}.$
We approximate the numerator and denominator separately. The construction of the approximations to the numerator and denominator is similar. In the following, we focus on the approximation of $q_{t}(\mathbf{x})$. The procedure for approximating $q_{t}(\mathbf{x})$ is outlined as follows:
Clipping the integral interval.
Recall that
$q_{t}(\mathbf{x})=\int_{\mathbb{R}^{d^{*}}}\widehat{p}_{data}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\underbrace{\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)}_{\text{Transition Kernel}}\mathrm{d}\mathbf{y}.$
According to Lemma F.2, there exists a constant $C>0$ such that for any $\mathbf{x}\in\mathbb{R}^{d^{*}}$,
$\left|q_{t}(\mathbf{x})-\int_{A_{\mathbf{x}}}\widehat{p}_{data}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)\mathrm{d}\mathbf{y}\right|\lesssim\epsilon,$
where $A_{\mathbf{x}}=\prod_{i=1}^{d^{*}}a_{i,\mathbf{x}}$ with $a_{i,\mathbf{x}}=\left[x_{i}-C\sigma_{t}\sqrt{\log\epsilon^{-1}},x_{i}+C\sigma_{t}\sqrt{\log\epsilon^{-1}}\right]$. This implies that we only need to approximate the integral over $A_{\mathbf{x}}$ sufficiently.
Approximating $\widehat{p}_{data}^{*}$.
Recall that the target density function $\widehat{p}_{data}^{*}(\mathbf{y})$ is assumed to be Hölder continuous, as stated in Assumption 4.4. Therefore, we can utilize the Taylor polynomial $f_{Taylor}^{*}(\mathbf{y})$ to approximate $\widehat{p}_{data}^{*}(\mathbf{y})$. Consequently, to approximate $q_{t}(\mathbf{x})$, we derive an approximation in the form of
$\int_{A_{\mathbf{x}}}f_{Taylor}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)\mathrm{d}\mathbf{y}.$
This analysis is detailed in Section B.1.
Approximating the transition kernel.
Although the Taylor polynomial $f_{Taylor}^{*}(\mathbf{y})$ can be implemented using a neural network, integrating over $\mathbf{y}$ remains challenging. To address this, we introduce the Taylor polynomial $f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})$ to approximate the exponential transition kernel, resulting in the following approximation:
$\int_{A_{\mathbf{x}}}f_{Taylor}^{*}(\mathbf{y})f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\mathrm{d}\mathbf{y}.$ (12)
See Section B.2 for more details. A similar approximation scheme using diffused local polynomials can also be applied to $\nabla q_{t}(\mathbf{x})$.
Approximating the integral via ReLU neural networks.
In (12), the product $f_{Taylor}^{*}(\mathbf{y})$ $\cdot f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})$ is a polynomial, thus its integration can be computed explicitly. Consequently, we use ReLU neural networks to approximate the diffused local polynomials, as outlined in Section B.3.
Finally, combining the above discussions, we derive the error bound for approximating $\nabla\log q_{t}(\mathbf{x})$ using a ReLU neural network, as detailed in Section B.4.
B.1 Approximating $\widehat{p}_{data}^{*}$ via local polynomials
In this section, we approximate $\widehat{p}_{data}^{*}$ via local polynomials, which is a key step for approximating the true score function $\nabla\log q_{t}(\mathbf{x})$. We give the following lemma.
Lemma B.1 (Approximating $\widehat{p}_{data}^{*}$ via local polynomials).
Suppose Assumptions 4.3-4.4 hold. Let $\mathcal{R}\gg 1$, there exists $p_{\mathcal{R}}(\mathbf{x})$ such that
$|\widehat{p}_{data}^{*}(\mathbf{x})-p_{\mathcal{R}}(\mathbf{x})|\lesssim\mathcal{R}^{-\beta},\leavevmode\nobreak\ \mathbf{x}\in[-1,1]^{d^{*}}.$
@@ B.1 Approximating $\widehat{p}_{data}^{*}$ via local polynomials
We denote
$f(\mathbf{x}):=\widehat{p}_{data}^{*}(2\mathbf{x}-1),\leavevmode\nobreak\ \mathbf{x}\in[0,1]^{d^{*}}.$
By Assumption 4.4, we know that $\|f\|_{\mathcal{H}^{\beta}([0,1]^{d^{*}})}\leq 2^{r}R$. For any $\mathbf{n}=(n_{1},n_{2},\cdots,n_{d^{*}})^{\top}\in[\mathcal{R}]^{d^{*}}:=\{0,1,\cdots,\mathcal{R}\}^{d^{*}}$, we define
$\psi_{\mathbf{n}}(\mathbf{x}):=\mathbf{I}_{\{\mathbf{x}\in\left(\frac{\mathbf{n}-1}{\mathcal{R}},\frac{\mathbf{n}}{\mathcal{R}}\right]\}}=\prod_{i=1}^{d^{*}}\mathbf{I}_{\{x_{i}\in\left(\frac{n_{i}-1}{\mathcal{R}},\frac{n_{i}}{\mathcal{R}}\right]\}},\leavevmode\nobreak\ \mathbf{x}=(x_{1},x_{2},\cdots,x_{d^{*}})^{\top}\in[0,1]^{d^{*}}.$
Note that the functions $\{\psi_{\mathbf{n}}\}_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}$ form a partition of unity of the domain $[0,1]^{d^{*}}$, i.e.,
$\sum_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}\psi_{\mathbf{n}}(\mathbf{x})\equiv 1,\leavevmode\nobreak\ \mathbf{x}\in[0,1]^{d^{*}}.$
Denote $p_{\mathbf{n},\boldsymbol{\alpha}}(\mathbf{x}):=\psi_{\mathbf{n}}(\mathbf{x})\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}$ and $c_{\mathbf{n},\boldsymbol{\alpha}}:=\partial^{\boldsymbol{\alpha}}f(\frac{\mathbf{n}}{\mathcal{R}})/\boldsymbol{\alpha}!$. Then $p_{\mathbf{n},\boldsymbol{\alpha}}(\mathbf{x})$ is supported on $\{\mathbf{x}\in[0,1]^{d^{*}}:\mathbf{x}\in\left(\frac{\mathbf{n}-1}{\mathcal{R}},\frac{\mathbf{n}}{\mathcal{R}}\right]\}$. We define
$p(\mathbf{x}):=\sum_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}\sum_{\|\boldsymbol{\alpha}\|_{1}\leq r}c_{\mathbf{n},\boldsymbol{\alpha}}p_{\mathbf{n},\boldsymbol{\alpha}}(\mathbf{x}),\leavevmode\nobreak\ \mathbf{x}\in[0,1]^{d^{*}}.$
Then, we have
$|f(\mathbf{x})-p(\mathbf{x})|$ $=\left|\sum_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}\psi_{\mathbf{n}}(\mathbf{x})f(\mathbf{x})-\sum_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}\psi_{\mathbf{n}}(\mathbf{x})\sum_{\|\boldsymbol{\alpha}\|_{1}\leq r}c_{\mathbf{n},\boldsymbol{\alpha}}\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}\right|$
$\leq\sum_{\mathbf{n}\in[\mathcal{R}]^{d^{*}}}\psi_{\mathbf{n}}(\mathbf{x})\left|f(\mathbf{x})-\sum_{\|\boldsymbol{\alpha}\|_{1}\leq r}c_{\mathbf{n},\boldsymbol{\alpha}}\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}\right|.$
Using Taylor expansion, there exists $\theta\in[0,1]$ such that
$f(\mathbf{x})=\sum_{\|\boldsymbol{\alpha}\|_{1}<r}c_{\mathbf{n},\boldsymbol{\alpha}}\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}+\sum_{\|\boldsymbol{\alpha}\|_{1}=r}\frac{\partial^{\boldsymbol{\alpha}}f\left((1-\theta)\frac{\mathbf{n}}{\mathcal{R}}+\theta\mathbf{x}\right)}{\boldsymbol{\alpha}!}\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}.$
Thus, we have
$\psi_{\mathbf{n}}(\mathbf{x})\left|f(\mathbf{x})-\sum_{\|\boldsymbol{\alpha}\|_{1}\leq r}c_{\mathbf{n},\boldsymbol{\alpha}}\left(\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right)^{\boldsymbol{\alpha}}\right|$ $\leq 2^{r}R\psi_{\mathbf{n}}(\mathbf{x})\sum_{\|\boldsymbol{\alpha}\|_{1}=r}\frac{1}{\boldsymbol{\alpha}!}\left\|\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right\|_{\infty}^{\|\boldsymbol{\alpha}\|_{1}}\theta^{\gamma}\left\|\mathbf{x}-\frac{\mathbf{n}}{\mathcal{R}}\right\|_{\infty}^{\gamma}$
$\leq 2^{r}R\psi_{\mathbf{n}}(\mathbf{x})\cdot\mathcal{R}^{-\beta}\sum_{\|\boldsymbol{\alpha}\|_{1}=r}\frac{1}{\boldsymbol{\alpha}!}$
$\leq 2^{r}R\psi_{\mathbf{n}}(\mathbf{x})\cdot\mathcal{R}^{-\beta}\cdot(d^{*})^{r},$
which implies that
$|f(\mathbf{x})-p(\mathbf{x})|\leq(2d^{*})^{r}R\cdot\mathcal{R}^{-\beta}\lesssim\mathcal{R}^{-\beta}.$
Finally, we define
$p_{\mathcal{R}}(\mathbf{x})=p\left(\frac{\mathbf{x}+1}{2}\right),\leavevmode\nobreak\ \mathbf{x}\in[-1,1]^{d^{*}},$
then we have
$|\widehat{p}_{data}^{*}(\mathbf{x})-p_{\mathcal{R}}(\mathbf{x})|=\left|f\left(\frac{\mathbf{x}+1}{2}\right)-p\left(\frac{\mathbf{x}+1}{2}\right)\right|\lesssim\mathcal{R}^{-\beta}.$
@@
In Lemma B.1, we assume that Assumptions 4.3-4.4 hold. By additionally incorporating Assumption 4.5, we can derive the following results.
Lemma B.2.
Suppose Assumptions 4.3-4.5 hold. Let $\mathcal{R}\gg 1$, there exists $p_{\mathcal{R}}(\mathbf{x})$ that satisfies